OpenStudy (anonymous):

how to solve 3 radical 75 plus 2 radical 5 and the solution? (operations with radical expressions)

OpenStudy (anonymous):



OpenStudy (anonymous):

hmm ok i have think, that's new, wait a sec haha

OpenStudy (anonymous):

haha its cool. its kind of a confusing topic lol

OpenStudy (anonymous):

it's not like this is it? 

OpenStudy (anonymous):

\[3\sqrt{75}+2\sqrt{5}\]

OpenStudy (anonymous):

not possible!

OpenStudy (anonymous):

already done!

OpenStudy (anonymous):

yea it is but its just confusing. algebra is not fun. at least to my stupid no-math brain lol

OpenStudy (anonymous):

wait, did you get it?

OpenStudy (anonymous):

is that how the assignment was given?

OpenStudy (anonymous):

wait, what? im kinda confused lol

OpenStudy (anonymous):

\[75=25\times 3\] so \[\sqrt{75}=\sqrt{25\times 3}=\sqrt{25}\times \sqrt{3}=5\sqrt{3}\]

OpenStudy (anonymous):

but you cannot combine it with \(2\sqrt{5}\) because they are not like terms, so your answer is probably \(15\sqrt{3}+2\sqrt{5}\)

OpenStudy (anonymous):

ah of course

OpenStudy (anonymous):

unless you meant to write the cubed root, as in \(\sqrt[3]{75}\) in which case you can do nothing

OpenStudy (anonymous):

no no i think the answer is just the above one before Bounty said "ah of course"

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks guys!